How many gallons of 25% alcohol solution and 40% alcohol solution must be mixed to get 9 gallons of 30% alcohol solution?
Here's what I tried which has to be wrong...
25x+40y=30
.40+25(9-x)=30(9) ??? stumped and confused from hours of this.
Thank you !

Answer by lwsshak3(11628)

How many gallons of 25% alcohol solution and 40% alcohol solution must be mixed to get 9 gallons of 30% alcohol solution?
***
let x=amt of 25% solution to mix
9-x=amt of 40% solution to mix
..
25%x+40%(9-x)=30%*9
.25x+3.6-.40x=2.7
.15x=0.90
x=6
9-x=3
amt of 25% solution to mix=6 gal
amt of 40% solution to mix=3 gal